Circular No. 6690

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997ct IN ANONYMOUS GALAXY J. Mueller reports her discovery of a supernova (mag ∼ 17) located

8′′.7 east and 4′′ south of the center of a galaxy at α = 14^{h}07^{m}24^{s}, δ = +70^{o}26′.1 (equinox 2000.0). SN 1997ct was found on a IV-N plate taken on June 29 with the 1.2-m Oschin Schmidt Telescope by Mueller and J. D. Mendenhall in the course of the second Palomar Sky Survey. No object appears at the position on the digitized sky survey or on a red sky survey plate taken on 1991 June 7. An uncalibrated spectrogram taken on July 1 by S. R. Kulkarni, J. C. Clemens, A. Sivaramakrishnan, and D. A. Frail with the Hale 5-m telescope (+ double spectrograph) shows the object to be a supernova.

GRB 970616 A. Udalski, Warsaw University Observatory, reports on behalf of the

OGLE-2 collaboration: “I-band CCD images of the entire XTE/IPN error box of the GRB 970616 field were obtained with the 1.3-m Warsaw telescope at the Las Campanas Observatory on June 22.400 and 23.403 UT. A comparison of images revealed the presence of a faint object, close to detection limit (I ∼ 21), with coordinates α = 1^{h}18^{m}28^{s}.00, δ = −5^{o}28′05′′.5 (equinox 2000.0) on the 180-s image from June 22, but not visible on the 180-s image taken on June 23. The object was not detected on additional, much deeper 900-s images taken on June 24.408 and 25.431; it is marginally visible on the 900-s image taken on June 30.416, so it must have faded by at least 1.5 mag during that time. The object is located ∼ 4′.5 away from the southern edge of the XTE/IPN error box, and hence it is not likely to be the afterglow of GRB 970616. This was the only fast fading object in the GRB 970616 field on our frames. Apparently the background of variable objects is high, and it may contribute significantly to false afterglow detections. Images

may be retrieved from ftp://sirius.astrouw.edu.pl/pub/udalski/grb
or ftp://astro.princeton.edu/bp/970616.”

NOVA SCORPII 1997 Photometry by A. C. Gilmore, obtained with the Mount John 0.6-m

f/16 reflector on June 27.506 UT: V = 12.82 ± 0.03, U − B = −0.22 ± 0.03, B − V = +0.57 ± 0.07, V − R = +1.61 ± 0.04, V − I = +1.47 ± 0.06 (comparison stars E790, E782, and E739 from Cousins’ E region).

1997 July 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green